OpenStudy (.whitedragon.):

A cone has a radius of 8 inches and a height of 12 inches. What is the exact volume of the cone? A. 32π in3 B. 96π in3 C. 256π in3 D. 768π in3

OpenStudy (anonymous):

The formula for the volume of a cone is 1/3 * base area * height

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\times8^{2}\times \pi \times12\]

OpenStudy (anonymous):

\[Volume of a \triangle is 1/3πr ^{2}h\] 1/3=0.33(π)(8)^2(12)=0.33(3.14)(64)(12)=795.85016 D is closest to the answer I got. 246(3.14)=803.84

OpenStudy (chunkymonkay):